| #1811 |
21 Nov 11:40 Log created
|
ryanque
|
Start experiment
|
normal
|
ISR
|
|
| #1812 |
24 Nov 15:00 Antenna VSWR
|
|
Antennas
|
normal
|
ISR
|
|
| #1813 |
24 Nov 19:30 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1814 |
24 Nov 20:50 ATR and TR calibration
|
|
Spark Gaps
|
normal
|
ISR
|
|
| #1815 |
24 Nov 20:55 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1816 |
24 Nov 21:00 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1817 |
24 Nov 21:28 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1818 |
24 Nov 21:35 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1819 |
25 Nov 00:00 ATR and TR calibration
|
|
Spark Gaps
|
normal
|
ISR
|
|
| #1820 |
25 Nov 00:00 Figures of noise, rtis and others
|
|
Plot
|
normal
|
ISR
|
|
| #1821 |
25 Nov 00:30 Imaging experiment turned off
|
|
Acquisition system
|
normal
|
ISR
|
|
| #1822 |
25 Nov 01:03 Ares PC stops
|
|
Acquisition system
|
normal
|
ISR
|
|
| #1823 |
25 Nov 01:05 Ares PC restored
|
|
Acquisition system
|
normal
|
ISR
|
|
| #1824 |
25 Nov 02:20 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1825 |
25 Nov 02:30 Transmitters enabled. EW drift channel D is attenuated.
|
|
Transmitters
|
normal
|
ISR
|
|
| #1826 |
25 Nov 02:40 TR change, EW drift channel D is OK
|
|
Transmitters
|
normal
|
ISR
|
|
| #1827 |
25 Nov 02:51 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1828 |
25 Nov 02:59 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1829 |
25 Nov 03:05 TR change, EW drift channel A is OK
|
|
Transmitters
|
normal
|
ISR
|
|
| #1830 |
25 Nov 03:25 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1831 |
25 Nov 03:30 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1832 |
25 Nov 04:05 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1833 |
25 Nov 04:13 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1834 |
25 Nov 04:25 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1835 |
25 Nov 04:33 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1836 |
25 Nov 04:40 Spark gaps adjusment
|
|
Spark Gaps
|
normal
|
ISR
|
|
| #1837 |
25 Nov 05:30 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1838 |
25 Nov 05:35 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1839 |
24 Nov 19:30 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR+Imaging
|
|
| #1840 |
25 Nov 07:05 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1841 |
25 Nov 07:10 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1842 |
25 Nov 07:27 Transmitters disabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1843 |
25 Nov 07:32 Transmitters enabled
|
|
Transmitters
|
normal
|
ISR
|
|
| #1844 |
25 Nov 08:01 Transmitters were down due to adjustments
|
|
Transmitters
|
normal
|
ISR
|
|
| #1846 |
25 Nov 09:18 Fall of the transmitters
|
I. Manay
|
Transmitters
|
normal
|
ISR
|
|
| #1847 |
25 Nov 09:40 Fall of the transmitters
|
I. Manay
|
Transmitters
|
normal
|
ISR
|
|
| #1849 |
25 Nov 12:04 Restart of the Faraday acquisition system
|
I. Manay
|
Acquisition system
|
normal
|
ISR
|
|
| #1850 |
25 Nov 19:30 Transmitter power measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1851 |
25 Nov 22:40 Calibration of ATr and Trs
|
J. Lavado
|
Start experiment
|
normal
|
ISR
|
|
| #1852 |
26 Nov 00:00 Figures of rtis, noises and others
|
P. Condor
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1853 |
26 Nov 02:13 Imaging Off
|
P. Condor
|
Acquisition system
|
normal
|
ISR
|
|
| #1854 |
26 Nov 02:30 Calibration of ATRs and TRs
|
J. Lavado
|
Antennas
|
normal
|
ISR
|
|
| #1855 |
26 Nov 07:20 Transmitter power measurements
|
|
Transmitters
|
normal
|
ISR+Imaging
|
|
| #1856 |
26 Nov 05:45 Restart Faraday
|
P. Condor
|
Acquisition system
|
normal
|
ISR
|
|
| #1857 |
26 Nov 06:45 Calibration of ATRs and Trs
|
J. Lavado
|
Antennas
|
normal
|
ISR
|
|
| #1858 |
26 Nov 07:28 Problem with Driver 1 screen
|
P. Condor
|
Transmitters
|
normal
|
ISR
|
|
| #1859 |
26 Nov 07:30 Batch of Faraday restart
|
P. Condor
|
Acquisition system
|
normal
|
ISR
|
|
| #1860 |
26 Nov 07:37 Restart Faraday Batch
|
P. Condor
|
Acquisition system
|
normal
|
ISR
|
|
| #1862 |
26 Nov 10:30 Spark Gaps Adjusment
|
|
Spark Gaps
|
normal
|
ISR
|
|
| #1863 |
26 Nov 15:15 Spark Gaps Adjusment
|
|
Spark Gaps
|
normal
|
ISR
|
|
| #1865 |
26 Nov 19:30 Transmitter power measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1866 |
27 Nov 00:20 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1867 |
27 Nov 07:08 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1871 |
27 Nov 15:40 Interferences at B channel of ew_drift experiment
|
I. Manay
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1872 |
27 Nov 16:12 Spark gaps adjustment
|
J. Lavado
|
Spark Gaps
|
normal
|
ISR
|
|
| #1873 |
27 Nov 17:08 End of the moon pass
|
I. Manay
|
General
|
normal
|
ISR
|
|
| #1874 |
27 Nov 17:45 Imaging experiment has started
|
I. Manay
|
Acquisition system
|
normal
|
ISR
|
|
| #1875 |
27 Nov 18:25 PC of ew_drift experiment was restarted
|
I. Manay
|
Acquisition system
|
normal
|
ISR
|
|
| #1876 |
27 Nov 19:15 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR+Imaging
|
|
| #1877 |
27 Nov 23:13 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1878 |
28 Nov 07:10 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1879 |
28 Nov 01:00 Imaging experiment turned off
|
AValdez
|
Acquisition system
|
normal
|
ISR+Imaging
|
|
| #1880 |
28 Nov 03:20am Fall of the transmitters
|
AValdez
|
Transmitters
|
normal
|
ISR+Imaging
|
|
| #1881 |
28 Nov 04:00am SPARK GAPS ADJUSTMENT
|
AValdez
|
Spark Gaps
|
normal
|
ISR
|
|
| #1882 |
27 Nov 23:59 Figures of rtis, noises and others
|
AValdez
|
Plot
|
normal
|
ISR
|
|
| #1845 |
25 Nov 09:00 Only EW Drift on
|
I. Manay
|
General
|
high
|
ISR
|
|
| #1848 |
25 Nov 11:30 Faraday on - ISR Complete
|
I. Manay
|
General
|
high
|
ISR
|
|
| #1861 |
24 Nov 19:30 Start of the experiment
|
I. Manay
|
Start experiment
|
high
|
ISR
|
|
| #1864 |
26 Nov 20:46 TX Down
|
|
Transmitters
|
high
|
ISR
|
|
| #1868 |
27 Nov 08:36 Transmitter OFF
|
P. Condor
|
Transmitters
|
high
|
ISR
|
|
| #1869 |
27 Nov 10:40 Experiment was restarted
|
I. Manay
|
General
|
high
|
ISR
|
|
| #1883 |
28 Nov 07:58 Faraday Te and Ti
|
ryanque
|
Process/Analysis Program
|
high
|
ISR
|
|
| #1884 |
28 Nov 08:05 Experiment finished
|
ryanque
|
General
|
high
|
ISR
|
|
| #2074 |
22 Jan 12:30 Data was uploaded to PUMA
|
I. Manay
|
General
|
high
|
ISR+Imaging
|
|